In a quantitative study, 4.624 grams of a compound containing carbon, hydrogen, and oxygen only yielded 7.210 g of CO2 and 2.656 g of H2O in a combustion apparatus. Determine the empirical formula of the compound.
1
Expert's answer
2020-11-14T13:59:04-0500

n(CO2)=n(C)=m/M=7.21/44=0.164 (mol)

m(C)=0.164*12=1.968(g)

n(H2O)=0.5n(H)=2.656/18=0.148 (mol)

m(H)=2*0.148=0.296(g)

m(O)=4.624-1.968-0.296=2.36(g)

n(H)=2*0.148=0.296(mol)

n(O)=2.36/16=0.148(mol)

n(C):n(H):n(O)=0.164:0.296:0.148=1.1:2:1=11:20:10.

Empirical formula is C11H20O10